OpenStudy (tatumlee):

find the range of the function f(x)=4x-1 for the domain {-1,0,1,2,3} how do I do this?

zepdrix (zepdrix):

You plug the domain values into the function, one at a time, and your result is included in the range. Example, here is the first one:\[\large\rm f(-1)=4(-1)-1\]\[\large\rm f(-1)=-5\]So our range looks like this so far: \(\large\rm\{-5,~~,~~,~~,~~\}\) Where the domain value of -1 corresponds to the range value of -5. Make sense? :o

OpenStudy (tatumlee):

so would the next one be 3?

zepdrix (zepdrix):

For the domain value x=0, Hmm no we don't get 3 for the range. Did you mean the one after that? :o

OpenStudy (tatumlee):

Yes I did, was there more solving to do on the other one?

zepdrix (zepdrix):

No, you just skipped a number :) I did -1, 0 comes after that, then 1, <- this is the one you did then 2, then 3

OpenStudy (tatumlee):

I thought I was doing the one for 0, I don't know how it became the next ones answer

zepdrix (zepdrix):

Oh you thought you were doing 0? Ok maybe a little math error then :) let's check.\[\large\rm f(0)=4(0)-1\]4 times 0, is 0,\[\large\rm f(0)=0-1\]So that's not giving us 3, right? :)

OpenStudy (tatumlee):

I think I might have subtracted 1 from 4 when I was supposed to multiply 4 by 0

zepdrix (zepdrix):

So what do you get for that next one? :)

zepdrix (zepdrix):

Looks like -1, ya? 0-1 = -1 \[\large\rm\{-5,-1,~~,~~,~~\}\] How bout when you plug x=1 into the function?

OpenStudy (tatumlee):

this time it would be three because 1 x 4 is 4-1= 3

zepdrix (zepdrix):

\[\large\rm Range:\{-5,-1,~3,~~,~~\}\]Ah there we go :)

zepdrix (zepdrix):

Almost done ^^ Just a couple more!

OpenStudy (tatumlee):

7,11?

zepdrix (zepdrix):

yay good job \c:/\[\large\rm Range:\{-5,-1,~3,~7,11\}\]

OpenStudy (tatumlee):

thank you!